What is the appropriate risk-adjusted discount rate for a project, given that RF = 3%, ERM = 8%, and the project’s beta is 1.54? a) 4.62% b) 9.32% c) 10.70% d) 15.32%
Answer #1

As per Capital Asset Pricing Model (CAPM),

The discount rate = RF + (beta*(ERM - RF)) = 3 +(1.54*(8-3)) = 10.7%

hence , the correct option is c) 10.70%